The time it takes to give a man a shampoo and haircut is normally distributed with mean 22 minutes and standard deviation 3 minutes. Customers are scheduled every 30 minutes.
(a)
What is the probability that a male customer will take longer than the allotted time? (Round your answer to 4 decimal places.)

Probability
(b)
If three male customers are scheduled sequentially on the half-hour, what is the probability that all three will be finished within their allotted half-hour times? (Round your answer to 4 decimal places.)

Answer by edjones(8007)

a)
z=(30-22)/3=2.6667
.0038 probability that a male customer will take longer than the allotted time
.
b)
1-.0038=.9962
.9962^3=.9886 the probability that all three will be finished within their allotted half-hour times.
.
Ed